(1+ tan1)(1+tan2)...........(1+tan45)=2^y


y=?????


 

Grade:

1 Answers

HARIKRISHNAN M
12 years ago

TAN45=1=TAN(1+44)

1=(TAN1+TAN44)/(1-TAN1TAN44)

1=TAN1+TAN44+TAN1TAN44

2=1+TAN1+TAN44+TAN1TAN44

I.E

2=(1+TAN1)(1+TAN44)

SIMILARLY

2=(1+TAN2)(1+TAN43)

.................................

................................

2=(1+TAN22)(1+TAN23)

MULTYPLING WE GET

(1+TAN1)(1+TAN2)...........(1+TAN44)=2^22

MULTIPLYING BY (1+TAN45)=2 we get

the question which is equaal to 2^23

so y=23

hope you like the answer